# Quantum Langevin Dynamics for Optimization

### v2

@@ 3.2 Generalization to the high-dimensional case
Assume that $V(x)\colon\mathbb{R}^{d}\to\mathbb{R}$ is an $d$-dimensional continuously differentiable convex function and $\nabla^{2}V(x)<+\infty$ for any $x$ in the domain, where $\nabla^{2}V=\sum_{j=1}^{d}\frac{\partial^{2}V}{\partial x_{j}^{2}}$ is the Laplacian of $V$. Let $x^{*}$ be the unique global minimum of $V(x)$. Then the solution $\rho(t)$ to Eq. (19), with approximation $\hbar\Omega\ll kT$, satisfies that
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kdT+O(e^{-\eta t}),$ (39)
if the initial density matrix $\rho(0)$ satisfies $\left\langle{H}\right\rangle_{0}-2kdT\geq 0$, where $\langle\cdot\rangle_{t}$ denotes the expectation value with respect to the density matrix $\rho(t)$, i.e., $\langle\cdot\rangle_{t}=\mathrm{Tr}(\cdot\rho(t))$.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot\left(2kdT-\langle x\cdot\nabla V+p^{2}/m\rangle_{t}\right)$
$\leq$ $\eta\cdot(2kdT-\left\langle{H}\right\rangle_{t}),$ (41)
@@
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kdT+O(e^{-\eta t}).$ (42)
@@ 3.3 Analytical solution for the quadratic potentials
$\frac{\mathrm{d}}{\mathrm{d}t}\langle V\rangle_{t}=$$\cancel{\langle\frac{\mathrm{d}}{\mathrm{d}t}V\rangle_{t}}+\frac{\mathrm{i}}{\hbar}\langle[H,V]\rangle_{t}+\left\langle-L^{\dagger}[L,V]+\left[L^{\dagger},V\right]L\right\rangle_{t}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle\frac{p^{2}}{2m}+\frac{1}{2}m\Omega^{2}x^{2}\right\rangle_{t}\hbar^{2}\left(\frac{\mu^{2}}{m}+\nu^{2}m\Omega^{2}\right)-\eta\langle m\Omega^{2}x^{2}+p^{2}/m\rangle_{t}.$ (45)
@@ 3.4 Convergence of QLD in the quasar-convex case
Assume that $V$ is a continuously differentiable function, which satisfies $\nabla^{2}V(x)<+\infty$ and $x\cdot\nabla V(x)\geq rV(x)\geq 0$, where $r\in(0,1)$ is a constant. Let $x^{*}$ be the unique global minimum of $V(x)$. Then the solution $\rho(t)$ to Eq. (14), with approximation $\hbar\Omega\ll kT$, satisfies that
$\left\langle{V(x)}\right\rangle_{t}-V(x^{*})\leq\frac{2}{r}kT+O(e^{-\eta rt}),$ (65)
if the initial density matrix $\rho(0)$ satisfies $\left\langle{H}\right\rangle_{0}-\frac{2}{r}kT\geq 0$.
@@
$\left\langle{V(x)}\right\rangle_{t}\leq\frac{2}{r}kT+O(e^{-\eta rt}).$ (66)
In this case, Eq. (35) still holds with approximation $\hbar\Omega\ll kT$ and assumption $\nabla^{2}V(x)<+\infty$. With the constraint $x\cdot\nabla V(x)\geq rV(x)\geq 0$, Eq. (35) becomes
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot\left(2kT-\langle x\cdot\nabla V+p^{2}/m\rangle_{t}\right)$
$\leq$ $\eta\cdot\left(2kT-\langle rV+p^{2}/m\rangle_{t}\right)$
$=$ $\eta\cdot\left(2kT-\left\langle{r(V(x)+\frac{p^{2}}{2m})+(1-\frac{r}{2})\frac{p^{2}}{m}}\right\rangle_{t}\right)$
$\leq$ $-\eta r\left\langle{H(x)}\right\rangle_{t}+2\eta kT.$ (67)
Let $h(t)=(\left\langle{H}\right\rangle_{t}-\frac{2}{r}kT)e^{\eta rt}$, then verify that $\frac{\mathrm{d}h(t)}{\mathrm{d}t}\leq 0$ for $\forall t>0$ by Eq. (67). Consequently, we get $h(t)\leq h(0)$,
$\left\langle{H(x)}\right\rangle_{t}-\frac{2}{r}kT\leq\left(\left\langle{H}\right\rangle_{0}-\frac{2}{r}kT\right)e^{-\eta rt}=O(e^{-\eta rt}).$ (68)
@@
$\left\langle{V(x)}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}\leq\frac{2}{r}kT+O(e^{-\eta rt}).$ (69)
@@
To achieve a desired accuracy $\varepsilon$, i.e., $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\varepsilon$ when $t\to+\infty$, we only need to set $kT=r\varepsilon/2$ and ensure that the approximation $\hbar\Omega\ll kT$ still holds.
@@ 4 Numerical results of time-independent QLD
In this section, we conduct numerical expertiments for time-independent QLD, i.e., all parameters remain constant throughout the evolution. In Section 4.1, we initially delve into the origin of dissipation, grounding our understanding in the intuitive concept of spontaneous emission. Section 4.1 also demonstrates that spontaneous emission can endow the system with the ability to transit to the ground state and achieve stability as the size of the heat bath increases. Section 4.2 primarily discusses the numerical methods employed to solving QLD. These technical details are essential for maintaining numerical stability during numerical computation. We then employ numerical experiments to validate the theoretical results concerning the quadratic potential from Section 3.3, as detailed in Section 4.3. In the experiments of the quadratic potential, we found that the numerical results are in alignment with the theoretical predictions. This consistency validates the precision of our numerical experiments. Subsequently, roles of each parameter in QLD are individually discussed in Section 4.4. We discover that the parameter $\eta$ influences the speed of convergence in QLD, while the magnitudes of $\hbar$ and $T$ impact the system’s ability to escape from local minima. Finally, Section 4.5 presents a comparison of the performance between the time-independent QLD and its classical counterpart, the Fokker-Planck-Smoluchowski equation. This comparison demonstrates the consistency of the classical and quantum thermal effects.
@@ 4.3 Quadratic potentials
Parameter Setting: Resolution $n=128$, domain $x\in[-1,1]$, space step $\Delta x=\frac{2}{n}$, time step $\Delta t=0.05\Delta x^{2}$, total evolution time $t_{f}=10$, $\hbar=2.1108$, $kT=165.8080~{}(T=200)$, characteristic damping rate $\eta=5$, $\Omega=223.2728$, $m=1$, initial state $\psi_{0}(x)=\sqrt{p(x)}$.
@@ 4.4 Roles of each parameter
In this part, we discuss the role of each parameter in QLD separately. In Eq. (14), parameters $\eta$, $T(kT)$, $\Omega$ play important roles in the behavior of the system. In order to analyze the roles of these parameters, we fix other parameters and change one of $\eta$, $T(kT)$, $\Omega$ each time.
@@ 5 Time-dependent QLD
Based on the conclusions drawn in Section 4.5, QLD does not demonstrate a significant advantage in situations where classical algorithms already perform well. Conversely, in landscapes where classical algorithms exhibit shortcomings, QLD shows only marginal superiority. This implies that we have not fully leveraged the advantages inherent in quantum physics. Consequently, we propose a time-dependent QLD to address this limitation.
@@ 5.1.1 Time-dependent $\eta(t)$
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}\leq\eta(t)\cdot\left(2kT-\langle H\rangle_{t}\right),$ (75)
@@ 5.1.2 Time-dependent $\hbar(t)$
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kT+O(e^{-\eta t}).$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}\leq$ $\frac{\mathrm{d}\ln\hbar}{\mathrm{d}t}\cdot\left\langle{p^{2}/m}\right\rangle_{t}+\eta\cdot\left(2kT-\langle V+p^{2}/(2m)\rangle_{t}\right).$ (81)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}\leq\eta\cdot\left(2kT-\langle V+p^{2}/(2m)\rangle_{t}\right),$ (82)
@@ 2. $f(T)\geq T$ for $\forall T>0$,
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kf(T(t))+O(e^{-\eta t}),$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}+\eta\cdot\left(\left\langle{H}\right\rangle_{t}-2kT(t)\right)\leq 0.$ (83)
@@
$\frac{\mathrm{d}h(t)}{\mathrm{d}t}=$ $\left(\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{H}\right\rangle_{t}-2k\frac{\mathrm{d}f}{\mathrm{d}T}\frac{\mathrm{d}T}{\mathrm{d}t}\right)e^{\eta t}+(\left\langle{H}\right\rangle_{t}-2kf(T(t)))\eta e^{\eta t}.$ (84)
@@
$\left(\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{H}\right\rangle_{t}-2k\frac{\mathrm{d}f}{\mathrm{d}T}\frac{\mathrm{d}T}{\mathrm{d}t}\right)+\eta(\left\langle{H}\right\rangle_{t}-2kf(T(t)))$ (86)
$\leq$ $\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}+\eta\cdot\left(\left\langle{H}\right\rangle_{t}-2kT(t)\right)\leq 0,$ (87)
@@
Then $h(t)\leq h(0)$ implies $\left\langle{H}\right\rangle_{t}-2kf(T(t))\leq O(e^{-\eta t})$. Using Eq. (23), we get $\left\langle{V}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}\leq 2kf(T(t))+O(e^{-\eta t})$.
@@ 5.6 Results of the comparison
In above experiments, we fix the domain of $x$ to $[-1,1]$ and all algorithms share the same total evolution time $t_{f}=10$ and same time step $\Delta t,s=0.05\cdot(\frac{2}{128})^{2}$. Besides, all quantum algorithms share same resolution $n=128$. Seven objective functions in Table 1 with different features are tested for each algorithm.

### v3

@@ 3.2 Generalization to the high-dimensional case
[…]
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\text{$\zeta$}kdT+O(e^{-\eta t}),$ (34)
if the initial density matrix $\rho(0)$ satisfies $\left\langle{H}\right\rangle_{0}-\text{$\zeta$}kdT\geq 0$, where $\zeta$ denotes the constant $2+O\left(\frac{\hbar\Omega}{kT}\right)$, and $\langle\cdot\rangle_{t}$ denotes the expectation value with respect to the density matrix $\rho(t)$, i.e., $\langle\cdot\rangle_{t}=\mathrm{Tr}(\cdot\rho(t))$.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot\left(kdT\left(2+O\left(\frac{\hbar\Omega}{kT}\right)\right)-\langle x\cdot\nabla V+p^{2}/m\rangle_{t}\right)$
$\leq$ $\eta\cdot(\zeta kdT-\left\langle{H}\right\rangle_{t}),$ (36)
@@
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\text{$\zeta$}kdT+O(e^{-\eta t}).$ (37)
@@ 3.3 Analytical solution for the quadratic potentials
$\frac{\mathrm{d}}{\mathrm{d}t}\langle V\rangle_{t}=$$\cancel{\left\langle\frac{\mathrm{d}}{\mathrm{d}t}V\right\rangle_{t}}+\frac{\mathrm{i}}{\hbar}\langle[H,V]\rangle_{t}+\left\langle-L^{\dagger}[L,V]+\left[L^{\dagger},V\right]L\right\rangle_{t}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle\frac{p^{2}}{2m}+\frac{1}{2}m\Omega^{2}x^{2}\right\rangle_{t}=\hbar^{2}\left(\frac{\mu^{2}}{m}+\nu^{2}m\Omega^{2}\right)-\eta\langle m\Omega^{2}x^{2}+p^{2}/m\rangle_{t}.$ (40)
@@ 3.4 Convergence of QLD in the quasar-convex case
[…]
$\left\langle{V(x)}\right\rangle_{t}-V(x^{*})\leq\frac{\text{$\zeta$}}{r}kT+O(e^{-\eta rt}),$ (60)
if the initial density matrix $\rho(0)$ satisfies $\left\langle{H}\right\rangle_{0}-\frac{\text{$\zeta$}}{r}kT\geq 0$, where $\zeta$ denotes the constant $2+O\left(\frac{\hbar\Omega}{kT}\right)$.
@@
$\left\langle{V(x)}\right\rangle_{t}\leq\frac{\text{$\zeta$}}{r}kT+O(e^{-\eta rt}).$ (61)
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot\left(\text{$\zeta$}kT-\langle x\cdot\nabla V+p^{2}/m\rangle_{t}\right)$
$\leq$ $\eta\cdot\left(\text{$\zeta$}kT-\langle rV+p^{2}/m\rangle_{t}\right)$
$=$ $\eta\cdot\left(\text{$\zeta$}kT-\left\langle{r(V(x)+\frac{p^{2}}{2m})+(1-\frac{r}{2})\frac{p^{2}}{m}}\right\rangle_{t}\right)$
$\leq$ $-\eta r\left\langle{H(x)}\right\rangle_{t}+\text{$\zeta$}\eta kT.$ (62)
Let $h(t)=(\left\langle{H}\right\rangle_{t}-\frac{\text{$\zeta$}}{r}kT)e^{\eta rt}$, then verify that $\frac{\mathrm{d}h(t)}{\mathrm{d}t}\leq 0$ for $\forall t>0$ by Eq. (62). Consequently, we get $h(t)\leq h(0)$,
$\left\langle{H(x)}\right\rangle_{t}-\frac{\text{$\zeta$}}{r}kT\leq\left(\left\langle{H}\right\rangle_{0}-\frac{\text{$\zeta$}}{r}kT\right)e^{-\eta rt}=O(e^{-\eta rt}).$ (63)
@@
$\left\langle{V(x)}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}\leq\frac{\text{$\zeta$}}{r}kT+O(e^{-\eta rt}).$ (64)
@@
To achieve a desired accuracy $\varepsilon$, i.e., $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\varepsilon$ when $t\to+\infty$, we only need to set $kT=r\varepsilon/\text{$\zeta$}$ and ensure that the approximation $\hbar\Omega\ll kT$ still holds.
@@ 4 Numerical results of time-independent QLD
In this section, we conduct numerical experiments for time-independent QLD, i.e., all parameters remain constant throughout the evolution. In Section 4.1, we initially delve into the origin of dissipation, grounding our understanding in the intuitive concept of spontaneous emission. Section 4.1 also demonstrates that spontaneous emission can endow the system with the ability to transit to the ground state and achieve stability as the size of the heat bath increases. Section 4.2 primarily discusses the numerical methods employed to solving QLD. These technical details are essential for maintaining numerical stability during numerical computation. We then employ numerical experiments to validate the theoretical results concerning the quadratic potential from Section 3.3, as detailed in Section 4.3. In the experiments of the quadratic potential, we found that the numerical results are in alignment with the theoretical predictions. This consistency validates the precision of our numerical experiments. Subsequently, roles of each parameter in QLD are individually discussed in Section 4.4. We discover that the parameter $\eta$ influences the speed of convergence in QLD, while the magnitudes of $\hbar$ and $T$ impact the system’s ability to escape from local minima. Finally, Section 4.5 presents a comparison of the performance between the time-independent QLD and its classical counterpart, the Fokker-Planck-Smoluchowski equation. This comparison demonstrates the consistency of the classical and quantum thermal effects.
@@ 4.3 Quadratic potentials
Parameter Setting: Resolution $n=128$, domain $x\in[-1,1]$, space step $\Delta x=\frac{2}{n}$, time step $\Delta t=0.05\Delta x^{2}$, total evolution time $t_{f}=10$, $\hbar=2.1108$, $kT=165.8080~{}(T=200)$, characteristic damping rate $\eta=5$, $\Omega=223.2728$, $m=1$, initial state $\psi_{0}(x)=\sqrt{p(x)}$.  [^5] We note here that the condition $\hbar\Omega\ll kT$ is not required in quadratic cases, because according to Section 3.3, the convergence of QLD in quadratic potentials is derived analytically without any assumptions.
@@ 4.4 Roles of each parameter
In this part, we discuss the role of each parameter in QLD separately. In Eq. (9), parameters $\eta$, $T(kT)$, $\hbar$ play important roles in the behavior of the system. In order to analyze the roles of these parameters, we fix other parameters and change one of $\eta$, $T(kT)$, $\hbar$ each time.
@@ 5 Time-dependent QLD
Based on the conclusions drawn in Section 4.5, QLD does not demonstrate a significant advantage in situations where classical algorithms already perform well. Conversely, in landscapes where classical algorithms exhibit shortcomings, QLD shows only marginal superiority. This implies that we have not fully leveraged the advantages inherent in quantum physics. Consequently, we propose a time-dependent QLD, where we manipulate the quantum and thermal effect more precisely in order to deal with non-convexity in objective functions.
@@ 5.1.1 Time-dependent $\eta(t)$
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}\leq\eta(t)\cdot\left(\text{$\zeta$}kT-\langle H\rangle_{t}\right),$ (70)
@@ 5.1.2 Time-dependent $\hbar(t)$
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\text{$\zeta$}kT+O(e^{-\eta t}).$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}\leq$ $\frac{\mathrm{d}\ln\hbar}{\mathrm{d}t}\cdot\left\langle{p^{2}/m}\right\rangle_{t}+\eta\cdot\left(\text{$\zeta$}kT-\langle V+p^{2}/(2m)\rangle_{t}\right).$ (76)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}\leq\eta\cdot\left(\text{$\zeta$}kT-\langle V+p^{2}/(2m)\rangle_{t}\right),$ (77)
@@ 2. $f(T)\geq T$ for $\forall T>0$,
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\text{$\zeta$}kf(T(t))+O(e^{-\eta t}),$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}+\eta\cdot\left(\left\langle{H}\right\rangle_{t}-\text{$\zeta$}kT(t)\right)\leq 0.$ (78)
@@
$\frac{\mathrm{d}h(t)}{\mathrm{d}t}=$ $\left(\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{H}\right\rangle_{t}-\text{$\zeta$}k\frac{\mathrm{d}f}{\mathrm{d}T}\frac{\mathrm{d}T}{\mathrm{d}t}\right)e^{\eta t}+(\left\langle{H}\right\rangle_{t}-\text{$\zeta$}kf(T(t)))\eta e^{\eta t}.$ (79)
@@
$\left(\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{H}\right\rangle_{t}-\text{$\zeta$}k\frac{\mathrm{d}f}{\mathrm{d}T}\frac{\mathrm{d}T}{\mathrm{d}t}\right)+\eta(\left\langle{H}\right\rangle_{t}-\text{$\zeta$}kf(T(t)))$ (81)
$\leq$ $\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}+\eta\cdot\left(\left\langle{H}\right\rangle_{t}-\text{$\zeta$}kT(t)\right)\leq 0,$ (82)
@@
Then $h(t)\leq h(0)$ implies $\left\langle{H}\right\rangle_{t}-\text{$\zeta$}kf(T(t))\leq O(e^{-\eta t})$. Using Eq. (33), we get $\left\langle{V}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}\leq\text{$\zeta$}kf(T(t))+O(e^{-\eta t})$.
@@ 5.6 Results of the comparison
In above experiments, we fix the domain of $x$ to $[-1,1]$ and all algorithms share the same total evolution time $t_{f}=10$ and the same time step $\Delta t$ and $s=0.05\cdot(\frac{2}{128})^{2}$. Besides, all quantum algorithms share same resolution $n=128$. Seven objective functions in Table 1 with different features are tested for each algorithm.
@@
6 Comparison of query complexity for different algorithms
In this section, we delve into the comparison of query complexity among various algorithms, i.e., the classical query complexity to local information and the quantum query complexity to the evaluation oracle [58, 16]. We assume that the quantum evaluation oracle $O_{V}$ allows us to query the function value of $V$ in a coherent way
6.1 Quantum query complexity for QLD
Theorem 4.
Given a Lindblad equation with form Eq. (9),
${O}\left(t\left(\hbar^{2}n^{2}+\|V\|_{\max}+\frac{\eta mkT}{\hbar^{2}}+\eta n\right)\frac{\log(t(\hbar^{2}n^{2}+\|V\|_{\max}+\frac{\eta mkT}{\hbar^{2}}+\eta n)/\epsilon)}{\log\log(t(\hbar^{2}n^{2}+\|V\|_{\max}+\frac{\eta mkT}{\hbar^{2}}+\eta n)/\epsilon)}\right)$
Proof.
By Definition 3, it is not difficult to give an $(\alpha_{E_{k}},a,\xi_{E_{k}})$-block-encoded matrix $U_{E_{k}}$ of $E_{k}$ and an $(\alpha_{V},a,\xi_{V})$-block-encoded matrix $U_{V}$ of $V$, where $\alpha_{E_{k}}$ is ${O}(\frac{\hbar^{2}}{\Delta x^{2}})={O}(\hbar^{2}n^{2})$ and $\alpha_{V}$ is ${O}(\|V\|_{\max})$. Here, $m$ is set to $1$ as in the previous numerical experiments. Then, Proposition 3 derives a $(\beta,a+b,\xi_{E_{k}}+\xi_{V})$-block-encoded matrix $\tilde{W}_{H}$ of $H$ with $\beta\geq|\alpha_{E_{k}}|+|\alpha_{V}|$. In addition, Proposition 3 also tells one query to $\tilde{W}_{H}$ can be implemented by one query to $U_{E_{k}}$ and one query to $U_{V}$.
Now, we have gotten a $(\beta,a+b,\xi_{E_{k}}+\xi_{V})$-block-encoded matrix $\tilde{W}_{H}$ of $H$. In our Lindblad equation Eq. (9), the unique $L$ is given by $L=\mu x+\mathrm{i}\nu p$. Presenting $L$ in the matrix form by finite difference method, we have $L=\mu\cdot\text{diag}(x_{1},\ldots,x_{n})+\nu\hbar\cdot\frac{\mathbf{D}}{\Delta x}$, where $\mathbf{D}$ is the differential matrix associated to the central difference method. Then, we can easily give an $(\alpha_{L},a,\xi_{L})$-block-encoded matrix $U_{L}$ of $L$, where $\alpha_{L}$ is at most ${O}(\mu+\frac{\nu\hbar}{\Delta x})={O}(\mu+{\nu\hbar}{n})$.
Using Proposition 4, $\|\mathcal{L}\|_{\text{be}}=\beta+\alpha_{L}^{2}$, where the smallest $\beta$ we can achieve is $|\alpha_{E_{k}}|+|\alpha_{V}|={O}(\hbar^{2}n^{2})+{O}(\|V\|_{\max})$, which gives
$\|\mathcal{L}\|_{\text{be}}$ $=O\left(\hbar^{2}n^{2}+\|V\|_{\max}+\mu^{2}+\hbar^{2}\nu^{2}n^{2}+2\hbar\mu\nu n\right)$ (101)
$=O(\hbar^{2}n^{2}+\|V\|_{\max}+\mu^{2}+\frac{\eta^{2}n^{2}}{4\mu^{2}}+\eta n).$ (102)
Following Eq. (24) and Eq. (25),
$\|\mathcal{L}\|_{\text{be}}=O\left(\hbar^{2}n^{2}+\|V\|_{\max}+\frac{\eta mkT}{\hbar^{2}}+\eta n\right).$
So, $O\left(t(\|\mathcal{L}\|_{\text{be}})\frac{\log(t(\|\mathcal{L}\|_{\text{be}})/\epsilon)}{\log\log(t(\|\mathcal{L}\|_{\text{be}})/\epsilon)}\right)$ queries to $\tilde{W}_{H}$ and $U_{L}$ are needed to simulate $e^{\mathcal{L}t}$ with error at most $\epsilon$. Recall that each query to $\tilde{W}_{H}$ can be implemented by one query to $U_{E_{k}}$ and one query to $U_{V}$, and one query to $U_{V}$ can be applied by one query to oracle $O_{V}$. So, $O\left(t(\|\mathcal{L}\|_{\text{be}})\frac{\log(t(\|\mathcal{L}\|_{\text{be}})/\epsilon)}{\log\log(t(\|\mathcal{L}\|_{\text{be}})/\epsilon)}\right)$ queries to $O_{V}$ are needed to simulate $e^{\mathcal{L}t}$ with error at most $\epsilon$. ∎
Definition 3 (Block-encoding, Definition 43 of [33]).
Definition 4 (State preparation pair, Definition 51 of [33]).
Suppose we are given an $(\alpha_{0},a,\epsilon^{\prime})$-block-encoding $U_{H}$ of $H$ and an $(\alpha_{i},a,\epsilon^{\prime})$-block-encoding $U_{L_{j}}$ for each $L_{j}$ $(j=1,\ldots,m)$. For all $t,\epsilon\geq 0$ with $\epsilon^{\prime}\leq\epsilon/(t\|\mathcal{L}\|_{\text{be}})$, there exists a quantum algorithm for simulating $e^{\mathcal{L}t}$ using
$O\left(t\|\mathcal{L}\|_{\text{be}}\frac{\log(t\|\mathcal{L}\|_{\text{be}}/\epsilon)}{\log\log(t\|\mathcal{L}\|_{\text{be}}/\epsilon)}\right)$
queries to $U_{H}$ and $U_{L_{j}}$, and some additional $1$- or $2$- qubit gates, where $\|\mathcal{L}\|_{\text{be}}=\alpha_{0}+\sum_{j=1}^{m}\alpha_{j}^{2}$.
6.2 Query complexity of other algorithms
6.2.1 Quantum query complexity for QHD
Because QHD simulates a time-dependent Hamiltonian system, its quantum query complexity can be inferred from the complexity involved in the simulation of the time-dependent Schrödinger equation. Here, we mainly follow Refs. [15, 53] and use quantum simulation in the real space to calculate the quantum query complexity for QHD in terms of the quantum oracle $O_{V}$ (Eq. (100)).
6.2.2 Quantum query complexity for QAA
6.2.3 Classical query complexity for SGD
6.2.4 Classical query complexity for NAGD
6.3 Comparison of query complexity for different algorithms
We also need to note that the roles of $t$ in QLD and QAA are different. Indeed, the interpretation of $t$ as mixing time in QLD, influenced by both the Hamiltonian and dissipative components, differs from the adiabatic scaling time in QAA, which relates to the gap of the time-dependent Hamiltonian. As a result, the time scales required for each algorithm to achieve optimal performance can differ significantly. In our numerical experiments, we only try to compare the performance of each algorithm at the same time scale $t$.